An R, L and C series circuit consists of a resistor of 100 ohms and an inductance of 100mh and a capacitance of 10 micro micro farad. If the voltage of 100v is applied across the combination. Find the resonance Q factor of the circuit.
1
Expert's answer
2019-03-21T16:16:20-0400
"Q=\frac{\rho}{R}=\sqrt{\frac{L}{C}}\cdot\frac{1}{R}=0.316."

"\rho" is called wave impedance, "Q" is quality factor.